# Recognising perfect fits

### v1

@@ 1   Introduction
Figure 1. The main goal of the report: to decide when we can pass between a veering triangulation (left) and a box decomposition for a pseudo-Anosov flow (right).
@@
Inspired by the rich computational theory for three-manifolds [haken-normal, matveev, kuperberg], we consider pseudo-Anosov flows from an algorithmic perspective. A natural problem is deciding when we can pass between a box decomposition and a veering triangulation; this is our main result. {restatable*}thmmainthm There is an algorithm that, given a box decomposition of a pseudo-Anosov flow, determines whether or not the flow has perfect fits. The case where the flow is Anosov is fundamentally distinct from the pseudo-Anosov case and is not covered by the theorem. Anosov flows are covered in a generalised version (5.1) which instead determines whether drilling orbits will produce a veering triangulation.
Given a decomposition for a three-manifold into flow boxes, deciding if this represents a pseudo-Anosov flow is straightforward (2.5), so Figure 1 addresses the main difficulty of the problem. Moreover, transitive pseudo-Anosov flows can be described by other discrete data, such as either veering triangulations equipped with appropriate filling slopes. This can be converted to a box decomposition ([agol-tsang, ss-pair]), to which we can apply the theorem.
@@
With the resulting veering triangulation, we make progress towards the natural recognition problem for flows. Recall that two pseudo-Anosov flows are orbit equivalent if there is a homeomorphism between the underlying manifolds that carries orbits to orbits. For a class of pseudo-Anosov flows, we say the orbit equivalence problem is solvable if, given box decompositions of two flows in the class, there exists an algorithm to decide if the flows are orbit equivalent. Strong results in the literature of [landry-minsky-taylor, bfm-classi, tsang-friedconjecture] imply that the data of the veering triangulation and the underlying closed manifold completely characterises the flow [tsang-friedconjecture, Corollary 6.2], giving us the following. {restatable*}coroequiv The orbit equivalence problem for pseudo-Anosov flows without perfect fits is solvable.
We can also use the veering triangulation to recognise if the underlying flow is a pseudo-Anosov suspension; this is 5.2.
@@
Acknowledgements. The author thanks Saul Schleimer for many interesting discussions about veering triangulations and pseudo-Anosov flows and for feedback on the paper. We also thank Chi Cheuk Tsang for many helpful comments on an early version of the paper.
@@ 2.1   Flows and box decompositions
For a hyperbolic matrix $A\in SL_{2}(\operatorname{\mathbb{Z}})$, the linear map $A:\operatorname{\mathbb{R}}^{2}\to\operatorname{\mathbb{R}}^{2}$ preserves a pair of foliations by lines on $\operatorname{\mathbb{R}}^{2}$. Consider $\operatorname{\mathbb{R}}^{2}$ as a copy of $\operatorname{\mathbb{C}}$. For $n\geq 2$, let $f_{k}:\operatorname{\mathbb{C}}\to\operatorname{\mathbb{C}}$ be a lift of $A$ under the semi-branched cover $z\to z^{k/2}$ which fixes the preimage of each quadrant of $\operatorname{\mathbb{R}}^{2}$. Define $f_{k,n}$ as $f_{k}$ followed by a rotation by $2\pi n/k$ for $0\leq n\leq k$. This map preserves the preimage of the horizontal and vertical foliations.
Suspend $f_{n,k}$ to obtain a flow $\psi_{n,k}$ on the resulting mapping torus of $\operatorname{\mathbb{R}}^{2}$. The map $\psi_{n,k}$ preserves the pair of two-dimensional foliations corresponding to the preimage of the invariant foliations of $A$. We denote the foliation whose leaves are contracted as $F^{s}$ and the other as $F^{u}$. See Figure 2(b).
@@
Let $M$ be a closed, oriented three-manifold. Let $\varphi$ be a continuous flow on $M$ with no stationary orbits. We call $(\varphi,\{s_{i}\})$ a pseudo-Anosov flow if there exists a pair of two-dimensional singular foliations denoted $M^{s}$ (stable) and $M^{u}$ (unstable) for which the following hold:
@@
It is conjectured ([mosher1996laminations, §3.1]) that the classes are equivalent for non-transitive flows. However, the notion of perfect fits can be defined for either variant of pseudo-Anosov flow, and a non-transitive always has perfect fits, see 2.13. ∎
@@ 2.2   Perfect fits
Let $M$ be a closed oriented three-manifold. Let $\varphi$ be a pseudo-Anosov flow on $M$ with $\{s_{i}\}$ the set of singular (or marked) orbits. The universal cover $\tilde{M}$ of $M$ is homeomorphic to $\operatorname{\mathbb{R}}^{3}$ (a consequence of [gabai-oertel, Theorem 6.1]). Lift the flow $\varphi$ to a flow $\tilde{\varphi}$ on $\tilde{M}$. The orbit space is the quotient $\mathcal{O}:=\tilde{M}/\tilde{\varphi}$. The orbit space is always homeomorphic to $\operatorname{\mathbb{R}}^{2}$ [fenley-mosher-qg, Proposition 4.1]. Moreover, $\mathcal{O}$ inherits a natural action of $\pi_{1}(M)$. The invariant foliations of $\tilde{\varphi}$ quotient to invariant singular stable and unstable foliations $\mathcal{O}^{u}$, $\mathcal{O}^{u}$ of $\mathcal{O}$.
@@
See Figure 3.
@@
We say that $\varphi$ has no genuine perfect fits if every perfect fit rectangle in $\mathcal{P}$ (or $\mathcal{O}^{\circ}$) is a cusp rectangle. This is called no perfect fits relative to $\{s_{i}\}$ in [agol-tsang, Definition 5.13]. ∎
Since singular orbits cannot appear on the interior of a lozenge, the generalised notion of perfect fits coincides with the classical notion for a pseudo-Anosov flow. However, the notions differ for a (pseudo-)Anosov flow with marked orbits (2.4).
@@ 2.3   Homotopic periodic orbits
Fenley characterises classical perfect fits by free homotopies. We say two periodic orbits $\alpha,\beta$ of $\varphi$ are almost freely homotopic if $\alpha^{n}\simeq\beta^{m}$. Following [fenley-annals, §4], we say that a periodic orbit is freely homotopic to itself if there exists a non-trivial element $\eta\in\pi_{1}(M)$ that conjugates $\alpha$ to $\alpha^{n}$ for some $n$. Either of these phenomena implies the existence of points in $\mathcal{O}$ that are fixed by the action of a common element in $\pi_{1}(M)$. In [fenley-chain, Theorem 3.3] and [fenley-good-geometry, Theorem 4.8], Fenley shows that if an element of $\pi_{1}(M)$ fixes distinct points in $\mathcal{O}$, there is a ‘chain of lozenges’ connecting the points in $\mathcal{O}$. In particular, such orbits imply that $\varphi$ has perfect fits. Fenley later used the shadowing lemma to prove the converse [fenley-qg, Theorem B].
@@
Let $\varphi$ be a pseudo-Anosov flow. Then $\varphi$ has perfect fits if and only if there is a pair of periodic points in $\mathcal{O}$ that are the corners of a lozenge. This equivalent to the existence of either:
@@
We will be more precise about the free homotopies. Suppose that periodic orbits $\alpha$ and $\beta$ periodic orbits (possibly the same) that form corners of a common lozenge in $\mathcal{P}$. So, we have $\alpha^{n}\simeq\beta^{m}$. Fenley shows in [fenley-almost-homotopic, Theorem 1.2] that when $\alpha$ and $\beta$ are regular orbits, $n$ and $m$ must have opposite signs, and we can take the powers $n$ and $m$ as having magnitude one or two. This is because the actions of the associated elements in $\pi_{1}(M)$ must fix the lozenge between the pair.
Suppose now that $\alpha$ and $\beta$ are singular. The generalisation is that $n$ and $m$ can be taken as having magnitudes the rotation periods (Item 2, 2.3) of $\beta$ and $\alpha$, respectively. ∎
For the generalised notion of genuine perfect fits, note that the foliation structure of a lozenge prohibits the presence of any singular orbits on its interior. So, the same criteria from above can be applied when the marked set of orbits consists precisely of singular orbits. In the case where $\{s_{i}\}$ contains regular orbits, we obtain a similar characterisation with some further analysis. First, we briefly recall generalised Dehn-Goodman-Fried surgery [goodman, fried-sections].
Construction 2.11 (Goodman, Fried).
@@
Let $(\varphi,\{s_{i}\})$ be a (pseudo-)Anosov flow with marked orbits. Let $\varphi(s)$ be a pseudo-Anosov flow with singular orbits filled into all of the marked orbits $\{s_{i}\}$. The two flows $\varphi$ and $\varphi(s)$ have the same lifted flow space $\mathcal{P}$, and one can observe (see [lmt-transverse, Remark 3.8.]) that $\varphi$ has genuine perfect fits if and only if $\varphi(s)$ has perfect fits in the classical sense. ∎
Orbits $\alpha$ and $\beta$ are almost freely homotopic when $\alpha^{n}$ and $\beta^{m}$ are conjugate in $\pi_{1}(M)$. The conjugacy problem is solvable in both the closed and with-boundary case by the work of [sela, Theorem 4.2] and Prèaux [preaux, Main Theorem]. These solutions are essential to both sides of the main algorithm.
@@
If a pseudo-Anosov flow $\varphi$ is non-transitive, the underlying manifold has an embedded, immersed transverse torus [mosher-transitive, Proposition 1.1]. Fenley shows that any pseudo-Anosov flow on a three-manifold $M$ with a $\operatorname{\mathbb{Z}}^{2}$-subgroup of $\pi_{1}(M)$ is either an Anosov suspension or has perfect fits [fenley-tori-give-fits, Main Theorem]. ∎
One could use 2.13 to immediately decide whether a flow has perfect fits by checking if the underlying manifold is toroidal and not a torus bundle. We do not explicitly need 2.13 for our algorithm, since the perfect fits are detected regardless. However, it highlights that detecting perfect fits is only interesting when $M$ is atoroidal. Pseudo-Anosov suspensions give examples of flows without perfect fits on atoroidal manifolds [fenley-cv-groups, Theorem G]. We are unaware of examples of pseudo-Anosov flows with perfect fits and singular orbits on atoroidal manifolds. We construct examples of such flows in 2.14.
@@ 2.4   Veering Triangulations
  • an assignment of a dihedral angle of 00 or $\pi$ to each edge of $t$ such that each vertex meets one $\pi$-angle edge and two 00-angle edges.
@@
  • Flow to triangulation: Let $\varphi$ be a pseudo-Anosov flow. If $\varphi$ has no genuine perfect fits, the drilled manifold $M^{\circ}$ admits a veering triangulation. This was announced by Agol-Guéritaud [ag-talk]; see [landry-minsky-taylor, §4] for an exposition. We use this explicitly in the form of 4.2.
@@
    Each filled torus has a periodic orbit of $\varphi$ at its core. With the filled cores included in the set of marked orbits, $\varphi$ has no genuine perfect fits. The precise prong data of the filled orbits are determined by the combinatorics of $\mathcal{V}$ on $\partial M^{\circ}$; we defer to the actual constructions Agol-Tsang [agol-tsang, Theorem 5.1] and Schleimer-Segerman [ss-pair, Theorem 10.1] for details. The construction of Schleimer-Segerman is an inverse to that of Agol-Guéritaud, which will be shown in an upcoming step of their program.
@@
Example 2.16.
Let $\mathcal{V}$ be a veering triangulation. Let $\varphi$ be a pseudo-Anosov flow obtained by filling regular periodic orbits into some of the cusps of $\varphi$. It is unclear under what conditions $\varphi$ has perfect fits in the classical sense. ∎
@@ 3   Finding a perfect fit
Let $\mathcal{B}$ be a box decomposition for a pseudo-Anosov flow $\varphi$. In this section, we give an algorithm (Algorithm 3.1) that will be used to verify if $\varphi$ has perfect fits.
@@ 3.1   Enumerating periodic orbits
For the second case, suppose that $\gamma$ lies inside the walls of $\mathcal{B}$, but not the one-skeleton. Choose a basepoint along $\gamma$ that lies in the interior of a wall $W$. There are either one or two push-offs $\hat{\gamma}_{i}$ of $\gamma$ into the interior of the boxes on either side of $W$. Since each $\hat{\gamma}_{i}$ is disjoint from the walls. They have an itinerary $\hat{w}_{i}$ as above. These are the itineraries of $\gamma$.
For the final case, suppose that $\gamma$ lies in the vertical one-skeleton of $\mathcal{B}$. Now $\gamma$ has wall push-offs into each of the walls whose boundaries meet at $\gamma$. These push-offs have itineraries as in the second case, which are the itineraries of $\gamma$. ∎
@@
Now suppose that there is a periodic orbit $\delta$ such that $\delta$ and $\gamma$ are corners of a common lozenge in $\mathcal{O}$. So, $\gamma^{i}\simeq\delta^{-j}$ where $i,j>0$ are the rotation periods of orbits 2.10. The preceding paragraph shows that the homotopy classes of $\gamma^{i}$ and $\delta^{j}$ are precisely the multiple encoded by their itineraries. ∎
Let $\mathcal{I}_{n}(\mathcal{B})$ be the set of itineraries with length at most $n$ that are lexically maximal among their own cyclic permutations. The hyperbolic dynamics of the flow $\varphi$ ensures that every symbolic orbit is realised: for $w\in\mathcal{I}_{n}(\mathcal{B})$, the corresponding loop $\gamma_{w}$ is homotopic to some multiple of a periodic orbit of $\varphi$.
@@
Observation 3.3.
By construction, duplicate itineraries only occur in the cases where an orbit $\gamma$ that meets the walls or lies in the vertical one-skeleton. Suppose first that $\gamma$ meets some wall $W$ of some box $B\in\mathcal{B}$. The Markov property implies that $\gamma$ is actually contained in a circular tuple of walls $(W_{i})$ such that each $W_{i+1}$ immediately above $W_{i}$. We call each wall $W_{i}$ a persistent wall and $(W_{i})$ a persistent tuple. See Figure 5. When $\gamma$ meets the vertical one-skeleton, the wall push-offs of $\gamma$ lie in a persistent tuple.
Figure 5. A persistent tuple of walls.
Starting just with $\mathcal{B}$, if $(W_{i})\subset\mathcal{B}$ is a persistent tuple, then there is a unique periodic orbit contained in the union of the $W_{i}$. Moreover, $(W_{i})$ has two push-offs, and these have itineraries from the second case of 3.1.
For $(W_{i}),(X_{j})$ two persistent tuples, define a relation by setting $(W_{i})\sim(X_{j})$ if $(W_{i})$ and $(X_{j})$ share an itinerary. The push-off equivalence is the transitive closure of $\sim$. This identifies duplicates: two itineraries in $\mathcal{I}_{n}(\mathcal{B})$ represent the same orbit if and only if they appear as itineraries for equivalent persistent tuples. ∎
@@
Let $S$ be an equivalence class of persistent tuples of walls. Each tuple in $S$ has two itineraries; we take the union over $S$ of all these itineraries and discard all but one from $\mathcal{I}_{n}(\mathcal{B})$. Call the resulting set ¯$\overline{\mathcal{I}}_{n}(\mathcal{B})$.
@@
Let $\gamma\subset M$ be a periodic orbit of $\varphi$ that has an itinerary of length at most $n$. Then there is a unique itinerary of $\gamma$ in ¯$\overline{\mathcal{I}}_{n}(\mathcal{B})$. ∎

### v2

@@ 1 Introduction
Figure 1. The main goal of the paper: to decide when we can pass between a veering triangulation (left) and a box decomposition for a pseudo-Anosov flow (right).
@@
Inspired by the rich computational theory for three-manifolds [haken-normal, matveev, kuperberg], we consider pseudo-Anosov flows from an algorithmic perspective. A natural problem is deciding when we can pass between a box decomposition and a veering triangulation. Our main result is solving this for pseudo-Anosov flows. The case of Anosov flows is fundamentally distinct from the pseudo-Anosov case, since there are not obvious orbits to drill. These are not addressed in this version of the theorem. {restatable*}thmmainthm There is an algorithm that, given a box decomposition of a pseudo-Anosov flow, determines whether or not the flow has perfect fits. Anosov flows are instead covered in a generalised version of the main theorem, (5.2), which determines whether drilling specified orbits will produce a veering triangulation.
Given a decomposition for a three-manifold into flow boxes, deciding if this represents a pseudo-Anosov flow is straightforward (2.5), so Figure 1 addresses the main difficulty of the problem. Moreover, transitive pseudo-Anosov flows can be described by other discrete data, such as a veering triangulation with appropriate filling slopes. This can be converted to a box decomposition ([agol-tsang, ss-pair]), to which we can apply the theorem.
@@
With the resulting veering triangulation, we make progress towards the natural recognition problem for flows. Recall that two pseudo-Anosov flows are orbit equivalent if there is a homeomorphism between the underlying manifolds that carries orbits to orbits. For a class of pseudo-Anosov flows, we say the orbit equivalence problem is solvable if, given box decompositions of two flows in the class, there exists an algorithm to decide if the flows are orbit equivalent.
In the absence of perfect fits, the veering triangulation and filling data are enough to characterise the flow [ss-final-step].
{restatable*}
coroequiv The orbit equivalence problem for pseudo-Anosov flows without perfect fits is solvable.
We can also use the veering triangulation to recognise if the underlying flow is a pseudo-Anosov suspension; this is 5.3.
@@
Acknowledgements. The author thanks Saul Schleimer for many interesting discussions about veering triangulations and pseudo-Anosov flows and for feedback on the paper. We thank Chi Cheuk Tsang for many helpful comments on an early version of the paper. Finally, we are very grateful to the anonymous reviewer; their detailed comments and corrections have greatly improved the paper.
@@ 2.1  Flows and box decompositions
[…]
@@
Let $M$ be a closed, oriented three-manifold. Let $\varphi$ be a continuous flow on $M$ with no stationary orbits. We call $\varphi$ a pseudo-Anosov flow if there exists a pair of two-dimensional singular foliations denoted $M^{s}$ (stable) and $M^{u}$ (unstable) for which the following hold:
@@
Given a pseudo-Anosov flow $\varphi$ and $\{s_{i}\}$ the set of singular orbits, we have the naturally associated pseudo-Anosov flow with marked orbits $(\varphi,\{s_{i}\})$. A box decomposition for $\varphi$ is also naturally one for $(\varphi,\{s_{i}\})$.
@@
It is conjectured ([mosher1996laminations, §3.1]) that the classes are equivalent for non-transitive flows. However, the notion of perfect fits can be defined for either variant of pseudo-Anosov flow, and a non-transitive always has perfect fits, see 2.14. ∎
@@ 2.2  Perfect fits
Let $M$ be a closed oriented three-manifold. Let $(\varphi,\{s_{i}\})$ be a (pseudo-)Anosov flow on $M$ with marked orbits. The universal cover $\tilde{M}$ of $M$ is homeomorphic to $\operatorname{\mathbb{R}}^{3}$ (a consequence of [gabai-oertel, Theorem 6.1]). Lift the flow $\varphi$ to a flow $\tilde{\varphi}$ on $\tilde{M}$. The orbit space is the quotient $\mathcal{O}:=\tilde{M}/\tilde{\varphi}$. The orbit space is always homeomorphic to $\operatorname{\mathbb{R}}^{2}$ [fenley-mosher-qg, Proposition 4.1]. Moreover, $\mathcal{O}$ inherits a natural action of $\pi_{1}(M)$. The invariant foliations of $\tilde{\varphi}$ quotient to invariant singular stable and unstable foliations $\mathcal{O}^{u}$, $\mathcal{O}^{u}$ of $\mathcal{O}$.
@@
A compact rectangle has boundary lying along two stable and two unstable leaves. See Figure 3.
@@
We say that $(\varphi,\{s_{i}\})$ has no genuine perfect fits if every perfect fit rectangle in $\mathcal{P}$ (or $\mathcal{O}^{\circ}$) is a cusp rectangle. This is called no perfect fits relative to $\{s_{i}\}$ in [agol-tsang, Definition 5.13]. ∎
Remark 2.9.
Let $\varphi$ be a pseudo-Anosov flow and $\{s_{i}\}$ the set of singular orbits. Since singular orbits cannot appear on the interior of a lozenge, then $\varphi$ has no perfect fits if and only if the pair $(\varphi,\{s_{i}\})$ has no genuine perfect fits. ∎
@@ 2.3  Homotopic periodic orbits
Fenley characterises classical perfect fits by free homotopies. We say two periodic orbits $\alpha,\beta$ of $\varphi$ are almost freely homotopic if $\alpha^{n}\simeq\beta^{m}$. Following [fenley-annals, §4], we say that a periodic orbit $\alpha$ is freely homotopic to itself if there exists a non-trivial element $\eta\in\pi_{1}(M)$, which itself is not a multiple of $\alpha$, that conjugates $\alpha$ to $\alpha^{n}$ for some $n$. Either of these phenomena implies the existence of points in $\mathcal{O}$ that are fixed by the action of a common element in $\pi_{1}(M)$. In [fenley-chain, Theorem 3.3] and [fenley-good-geometry, Theorem 4.8], Fenley shows that if an element of $\pi_{1}(M)$ fixes distinct points in $\mathcal{O}$, there is a ‘chain of lozenges’ connecting the points in $\mathcal{O}$. In particular, such orbits imply that $\varphi$ has perfect fits. Fenley later used the shadowing lemma to prove the converse [fenley-qg, Theorem B].
@@
Let $\varphi$ be a pseudo-Anosov flow. Then $\varphi$ has perfect fits if and only if there is a pair of periodic points in $\mathcal{O}$ that are the corners of a lozenge. This is equivalent to the existence of either:
@@
We will be more precise about the free homotopies. Suppose that $\alpha$ and $\beta$ are periodic orbits (possibly the same) that form corners of a common lozenge in $\mathcal{P}$. So, we have $\alpha^{n}\simeq\beta^{m}$. Fenley shows in [fenley-almost-homotopic, Theorem 1.2] that when $\alpha$ and $\beta$ are regular orbits, $n$ and $m$ must have opposite signs, and we can take the powers $n$ and $m$ as having magnitude one or two. This is because the actions of the associated elements in $\pi_{1}(M)$ must fix the lozenge between the pair.
Suppose now that $\alpha$ and $\beta$ are singular. The generalisation is that $n$ and $m$ can be taken as having magnitudes the rotation periods (Item 2, 2.3) of the orbits $\alpha$ and $\beta$, respectively. ∎
We can obtain a similar characterisation of no genuine perfect fits for flows with marked orbits. First, we briefly recall generalised Dehn-Goodman-Fried surgery [goodman, fried-sections].
Construction 2.12 (Goodman, Fried).
@@
Let $(\varphi,\{s_{i}\})$ be a (pseudo-)Anosov flow with marked orbits. Let $\varphi(s)$ be a pseudo-Anosov flow with singular orbits filled into all of the marked orbits $\{s_{i}\}$. The two flows $\varphi$ and $\varphi(s)$ have the same lifted flow space $\mathcal{P}$, and one can observe (see [lmt-transverse, Remark 3.8.]) that $(\varphi,\{s_{i}\})$ has genuine perfect fits if and only if $\varphi(s)$ has perfect fits. ∎
Orbits $\alpha$ and $\beta$ are almost freely homotopic when $\alpha^{n}$ and $\beta^{m}$ are conjugate in $\pi_{1}(M)$. The integers $m$ and $n$ can be negative. The conjugacy problem is solvable in both the closed and with-boundary case by work of [sela, Theorem 4.2] and Prèaux [preaux, Main Theorem]. These solutions are essential to both sides of the main algorithm.
@@
If a pseudo-Anosov flow $\varphi$ is non-transitive, the underlying manifold has an embedded transverse torus [mosher-transitive, Proposition 1.1]. Fenley shows that any pseudo-Anosov flow on a three-manifold $M$ with a $\operatorname{\mathbb{Z}}^{2}$-subgroup of $\pi_{1}(M)$ is either an Anosov suspension or has perfect fits [fenley-tori-give-fits, Main Theorem]. ∎
One could use 2.14 to immediately decide whether a flow has perfect fits by checking if the underlying manifold is toroidal and not a torus bundle. We do not explicitly need 2.14 for our algorithm, since the perfect fits are detected regardless. However, it highlights that detecting perfect fits is only interesting when $M$ is atoroidal. Pseudo-Anosov suspensions give examples of flows without perfect fits on atoroidal manifolds [fenley-cv-groups, Theorem G]. We give some simple examples of pseudo-Anosov flows with perfect fits and singular orbits on atoroidal manifolds. Another way to construct such examples is by using the ideas presented in [mosher-homology-II, Comment 2, §4].
@@ 2.4  Veering Triangulations
  • an assignment of a dihedral angle of 0 or $\pi$ to each edge of $t$ such that each vertex meets one $\pi$-angle edge and two 0-angle edges. Moreover, an edge between two faces with the same co-orientation must have dihedral angle $\pi$.
@@
  • Flow to triangulation: Let $\varphi$ be a pseudo-Anosov flow. If $\varphi$ has no perfect fits, the drilled manifold $M^{\circ}$ admits a veering triangulation. This was announced by Agol-Guéritaud [ag-talk]; see [landry-minsky-taylor, §4] for an exposition. We use this explicitly in the form of 4.2.
@@
    Each filled torus has a periodic orbit of $\varphi$ at its core. With the filled cores included in the set of marked orbits, $\varphi$ has no perfect fits. The precise prong data of the filled orbits are determined by the combinatorics of $\mathcal{V}$ on $\partial M^{\circ}$; we defer to the actual constructions Agol-Tsang [agol-tsang, Theorem 5.1] and Schleimer-Segerman [ss-pair, Theorem 10.1] for details. The construction of Schleimer-Segerman is an inverse to that of Agol-Guéritaud, which will be shown in an upcoming step of their program.
@@ 3 Finding a perfect fit
Let $\mathcal{B}$ be a box decomposition for a pseudo-Anosov flow $\varphi$. In this section, we give an algorithm (Algorithm 3.6) that will be used to verify if $\varphi$ has perfect fits. We adapt this to the setting of a (pseudo-)Anosov flow with marked orbits and genuine perfect fits in 3.8.
@@ 3.1  Enumerating periodic orbits
Now suppose that $\gamma$ meets the two-skeleton. The Markov property implies that $\gamma$ is contained in the union of a circular tuple of walls $\cup W_{i}$ such that each $W_{i+1}$ lies immediately above $W_{i}$. We call each wall $W_{i}$ a persistent wall and $(W_{i})$ a persistent tuple. See Figure 5.
Figure 5. A persistent tuple of walls.
Restrict further to when $\gamma$ meets the walls of $\mathcal{B}$, but not the vertical one-skeleton. This means that $\gamma$ is non-singular, and $\gamma$ is contained in the union of a persistent tuple $(W_{i})$. The intersection $x=\gamma\cap W_{i}$ is a one-cell. We can push $x$ into the box on either side of $W_{i}$ to obtain a vertical segment disjoint from the walls. The push-off moves each endpoint of $x$ from the boundary of a rectangle to its interior. We take both choices of push-off, and repeat for all the $W_{i}$. Since the pushed-off segments for $W_{i}$ and $W_{i+1}$ naturally glue top to bottom, all of the segments stitch together to give either one or two curves $\hat{\gamma}_{i}$. We call these the push-offs of $\gamma$. Since each $\hat{\gamma}_{i}$ is disjoint from the walls, they have an itinerary $\hat{w}_{i}$ as above. These are the itineraries of $\gamma$.
For the remaining case, suppose that $\gamma$ lies in the vertical one-skeleton of $\mathcal{B}$. Now $\gamma$ is a union of vertical one-cells. Such a one-cell $x$ meets the boundary of multiple persistent walls. We can then push-off $x$ into any of these adjacent walls. We do so for each such wall, and repeat for all choices of $x$. The resulting vertical segments stitch together to give a collection of loops $\{\hat{\gamma}_{i}\}$. The precise number of these loops depends on the prong-type and rotation period of $\gamma$. Each $\hat{\gamma}_{i}$ is contained in the union of a persistent tuple, so we can apply the previous case to obtain a set of itineraries. Taking the union of these itineraries over all $\hat{\gamma}_{i}$ gives the itineraries of $\gamma$. ∎
@@
Now suppose that there is a periodic orbit $\delta$ such that $\delta$ and $\gamma$ are corners of a common lozenge in $\mathcal{O}$. So, $\gamma^{i}\simeq\delta^{-j}$ where $i,j>0$ are the rotation periods of orbits 2.11. The preceding paragraph shows that the homotopy classes of $\gamma^{i}$ and $\delta^{j}$ are precisely the multiple encoded by their itineraries. ∎
Choose an ordering on the top/bottom rectangles of the boxes. This endows itineraries with a lexical order. Let $\mathcal{I}_{n}(\mathcal{B})$ be the set of itineraries with length at most $n$ that are lexically maximal among their own cyclic permutations.
Lemma 3.3.
Two distinct periodic orbits $\alpha$ and $\beta$ of $\varphi$ have distinct itineraries in $\mathcal{I}_{n}(\mathcal{B})$.
Proof.
Let $w\in\mathcal{I}_{n}(\mathcal{B})$ be a shared itinerary of $\alpha$ and $\beta$. For $r$ a subrectangle on the bottom of box $B$, let $f(r)$ be the image of $r$ in the top of $B$ obtained by flowing vertically upwards. Let $(r_{i})$ be the infinite sequence of rectangles passed through by following $w$ periodically forward. Set $s_{0}=r_{0}$ and $s_{i+1}=f(s_{i})\cap r_{i+1}$. Then $\alpha$ and $\beta$ intersect $s_{i}$, while the Markov property forces $s_{i}$ to contract to a stable segment as $i\to\infty$. This implies $\alpha=\beta$. ∎
@@
Observation 3.4.
By construction, duplicate itineraries only occur in the cases where an orbit $\gamma$ that meets the walls or lies in the vertical one-skeleton.
Starting just with $\mathcal{B}$, if $(W_{i})\subset\mathcal{B}$ is a persistent tuple, then there is a unique periodic orbit contained in the union of the $W_{i}$. The itineraries of this orbit are those of its push-offs as in 3.1.
For $(W_{i}),(X_{j})$ two persistent tuples, define a relation by setting $(W_{i})\sim(X_{j})$ if the orbits covered by $(W_{i})$ and $(X_{j})$ share an itinerary. The push-off equivalence is the transitive closure of $\sim$. Because of Lemma 3.3, this identifies all duplicates: two itineraries in $\mathcal{I}_{n}(\mathcal{B})$ represent the same orbit if and only if they appear as itineraries for equivalent persistent tuples. ∎
@@
Let $S$ be an equivalence class of persistent tuples of walls. Each tuple in $S$ has two itineraries; we take the union over $S$ of all these itineraries and discard all but one from $\mathcal{I}_{n}(\mathcal{B})$. Call the resulting set $\overline{\mathcal{I}}_{n}(\mathcal{B})$.
@@
Let $\gamma\subset M$ be a periodic orbit of $\varphi$ that has an itinerary of length at most $n$. Then there is a unique itinerary of $\gamma$ in $\overline{\mathcal{I}}_{n}(\mathcal{B})$. ∎